OpenStudy (anonymous):

If X is 2/3 of Y and Y is 3/4 Of Z, what is the ratio of Z:X? Here is what I did: X=2/3Y Y=3/4Z X=2/3(3/4Z) substituted Y X=1/2 Z multiplied the fractions 2x=1z multiplied by 2 so z:x=1:2 but apparently the right answer is 2:1, so what did I do wrong?

OpenStudy (anonymous):

you are correct upto x= 1/2z divide all through by z this gives you: x/z = 1/2 or x:z = 1:2 or z:x = 2:1

OpenStudy (anonymous):

you lost me at divide all through by z.

OpenStudy (anonymous):

see, u got 2x = 1z right? now just divide throughout by x. what do u get?

OpenStudy (anonymous):

so i should divide by x? then it's 2=z/x?

OpenStudy (anonymous):

whatever you divide with, you'll get the right answer, so if you decide to divide by x then you have 2 =z/x, but isn't two the same as 2/1

OpenStudy (anonymous):

exactly! 2 = z/x or 2/1 = z/x or 2:1 = z:x got it?

OpenStudy (anonymous):

nope I still don't get it why would you divide by the variable anyway?

OpenStudy (anonymous):

u want x in the denominator right?

OpenStudy (anonymous):

2x=1z is a ration doesn't that mean for every 2Xs there's 1Z so the ratio of z:x is 1:2

OpenStudy (anonymous):

what is wrong with what I just did?

OpenStudy (anonymous):

are u confusing a ratio with the proportional sign? see, a ratio means something divided by something. so z:x = z/x. is this clear?

OpenStudy (anonymous):

2x=1z is correct, but remember you need a ratio, so you have to put like terms together. in your case (x,z) is one set of like terms and (2,1) is another, so what do you do to this equation to bring like terms to the same side? can you please cross multiply x/z =1/2 and see what you get

OpenStudy (anonymous):

2x=1z

OpenStudy (anonymous):

ok i got it now, I have to get the variables together to make it right

OpenStudy (anonymous):

thanks both of you

OpenStudy (anonymous):

np^^